I'm having trouble solving using the recursive formula. Here's the problem:
"Find the first five terms of the sequence given the following using the recursive formula-
a-sub-1 = 8 and a-sub-n = 10-3a-sub-n minus 1, for n ≥ 2

Answer by Edwin McCravy(20056)

a1 = 8, an = 10-3an-1, for n ≥ 2
Substitute n=2 in an = 10-3an-1

an = 10-3an-1

a2 = 10-3a2-1

a2 = 10-3a1

Substitute 8 for a1

a2 = 10-3(8)

a2 = 10-24

a2 = -14

-------------------------------

Substitute n=3 in an = 10-3an-1

an = 10-3an-1

a3 = 10-3a3-1

a3 = 10-3a2

Substitute -14 for a2

a3 = 10-3(-14)

a3 = 10+42

a3 = 52

-------------------------------

Substitute n=4 in an = 10-3an-1

an = 10-3an-1

a4 = 10-3a4-1

a4 = 10-3a3

Substitute 52 for a3

a4 = 10-3(52)

a4 = 10-156

a4 = -146

-------------------------------

Substitute n=5 in an = 10-3an-1

an = 10-3an-1

a5 = 10-3a5-1

a5 = 10-3a4

Substitute -146 for a4

a5 = 10-3(-146)

a5 = 10+438

a5 = 448

-------------------------------

Edwin